# Yintong Huo‡§, Michael R. Lyu

### v1

@@ V-B Metric
Visual Metrics. CLIP [48] is applied to measure the semantic similarity between the generated and original webpages.
@@
Code Metrics. (1) Compilation Success Rate (CSR). This metric represents the percentage of generated code that compiles successfully without errors. Assume that the total number of samples is N and the number of samples compiled successfully is S, then $CSR=\frac{S}{N}$. (2) Code Modification Similarity (CMS). We employ the Jaccard similarity [49] to quantify the precision of code modifications on design edit and design repair tasks by comparing the sets of modified line numbers between the ground truth and generated code. Let $A$ represent the set of line numbers modified in the ground truth code and $B$ represent the set of line numbers modified in the generated code. The CMS is formally defined as: $CMS(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
@@
MLLM-as-Judge Metrics. MLLMs have shown great performance in assisting judges across diverse modalities [50, 51]. Therefore, we prompt GPT-4o [42] to determine whether the model meets the user’s requirements on the design edit task and resolve the design issues on the design repair task, and output an MLLM score between 0 and 10 with detailed explanations (0-3 denotes the poor edit/repair, 4-6 denotes partial edit/repair, 7-8 denotes Good eidt/repair and 9-10 denotes excellent edit/repair). For design edit task, we sample 359 samples and validate this MLLM score through human evaluation, which can achieve an average accuracy of 95.54%. For design repair task, we sample 111 samples and validate this MLLM score through human evaluation, which can achieve an average accuracy of 91.89%.
TABLE III: Studied Multimodal Large Language Models.
@@
Pixtral [43] Pixtral-12B-2409 Pixtral-12B 12B
Pixtral [43] Pixtral-large-latest Pixtral-124B 124B
Qwen [44] Qwen2.5-VL-7B-Instruct Qwen-7B 7B
Qwen [44] Qwen2.5-VL-72B-Instruct Qwen-72B 72B
LLama [52] Llama-3.2-11B-Vision-Instruct Llama-11B 11B
LLama [52] Llama-3.2-90B-Vision-Instruct LLama-90B 90B
Gemini [46] Gemini-2.0-Flash Gemini-2.0 -
GPT [42] GPT-4o-2024-11-20 GPT-4o -
Claude [47] Claude-3-7-sonnet-20250219 Claude-3.7 -
V-C Research Questions
@@
  • RQ3: (Influence of difficulty) How does varying task difficulty impact MLLM performance?
  • RQ4: (Influence of context) To what extent do different input contexts affect MLLMs’ performance?
@@
VI Experiment Results
TABLE IV: The model performance on DesignBench under different tasks and different frameworks. Bold numbers on a dark red background indicate the maximum values, and an underline with a light red background denotes the second-best value.
@@ VI Experiment Results
Compilation React 0.9541 0.9725 0.9083 0.9450 0.8991 0.9725 0.8532 0.9541 0.1284
@@
CMS Vue 0.4050 0.3698 0.3276 0.2104 0.0655 0.3046 0.2394 0.3276 0.0862
React 0.4659 0.5246 0.3710 0.2637 0.1819 0.4093 0.4120 0.4398 0.0815
Angular 0.6829 0.6099 0.6392 0.4700 0.3621 0.3264 0.2867 0.6018 0.1367
Vanilla 0.3439 0.3394 0.2905 0.1946 0.1582 0.3651 0.2770 0.3209 0.1635
Compilation React 1.0000 0.9815 1.0000 0.9167 0.7963 0.9907 1.0000 0.9907 0.4815
@@
CMS React 0.4827 0.2752 0.1755 0.0448 0.0473 0.2272 0.0905 0.1866 0.0417
Vue 0.3065 0.2524 0.1782 0.0501 0.0474 0.2230 0.1557 0.1131 0.0127
Angular 0.5719 0.5073 0.3968 0.3099 0.2688 0.2618 0.2546 0.5563 0.0000
Vanilla 0.2287 0.1637 0.1630 0.0365 0.0690 0.1398 0.1188 0.1446 0.0277
Compilation React 1.0000 1.0000 1.0000 0.9286 0.9643 1.0000 0.9643 0.9286 0.2857
@@
VI-A RQ1: How do MLLMs perform across distinct front-end tasks?
Table IV presents the performance of nine MLLMs on three tasks in three front-end frameworks React, Vue, and Angular. Vanilla denotes the webpage developed by vanilla HTML/CSS.
@@ VI-A RQ1: How do MLLMs perform across distinct front-end tasks?
In Design Generation tasks, MLLMs face two primary bottlenecks: compilation errors and visual inaccuracies. The compilation rates reveal significant framework-dependent challenges, with Angular showing the lowest success rates (0.6867-0.7590 for top models) compared to React and Vue ($>$0.83 for top models). Additionally, the moderate CLIP scores (0.5963-0.6324 even for best performers) indicate substantial room for improvement in generating visually accurate webpage layouts. The performance gap between vanilla HTML (highest CLIP scores) and framework-based implementations further suggests that the syntactic complexity in modern frameworks exacerbates challenges in both compilation and visual rendering.
In Design Edit and Design Repair tasks, the primary bottleneck of MLLMs is the accurate localization of code segments requiring modification. This is evidenced by CMS scores that are lower than compilation rates across all models. In Design Edit, even top-performing models like Claude-3.7 achieve CMS scores of only 0.3439-0.6829, despite maintaining compilation rates above 0.9. Similarly, in Design Repair, CMS scores range from 0.2287-0.5719 for Claude-3.7. These results highlight a substantial difficulty in localizing target code, even when the generated code successfully compiles.
Finding 3: MLLMs exhibit task-specific bottlenecks: Design Generation is challenged by compilation errors and visual inaccuracies, while Design Edit and Repair are mainly limited by deficiencies in code localization.
VI-B RQ2: What is the comparative performance of MLLMs when applied to different development frameworks?
@@
Figure 6: Performance of webpages implemented by different frameworks. y-axis denotes the framework used to actually implement the webpage and x-axis represents the framework used by the model.
TABLE V: Performance under different difficulty levels.
Model Design Generation Design Edit Design Repair
Easy Medium Hard Easy Medium Hard Easy Medium Hard
Claude-3.7 0.83 0.76 0.73 8.64 8.32 7.61 7.21 6.81 6.93
GPT-4o 0.79 0.74 0.69 8.73 7.95 8.19 6.93 6.46 5.53
Gemini-2.0 0.80 0.73 0.70 9.05 7.84 7.51 7.07 6.21 4.19
Llama-90B 0.68 0.64 0.62 6.54 6.83 5.68 5.35 4.90 4.39
Llama-11B 0.65 0.60 0.59 5.88 5.05 5.24 3.91 4.07 2.41
Pixtral-124B 0.78 0.72 0.71 8.69 7.91 6.98 7.30 6.56 5.66
Pixtral-12B 0.70 0.66 0.68 8.26 7.63 6.71 6.76 5.84 5.30
Qwen-72B 0.76 0.69 0.69 7.78 6.81 6.72 7.12 6.06 5.57
Qwen-7B 0.55 0.53 0.52 3.46 2.83 2.71 1.09 1.63 1.05
@@
VI-C RQ3: How does varying task difficulty impact MLLM performance?
We categorize samples into different difficulty levels to systematically evaluate MLLMs’ performance across varying complexity scenarios. The difficulty assessment criteria differ for each task to reflect their distinct challenges.
For Design Generation task, which require generating code from scratch, difficulty is primarily determined by the visual complexity of webpages. We classify webpages smaller than 1000$\times$1000 pixels as easy, those larger than 2000$\times$2000 pixels as hard, and intermediate sizes as medium difficulty.
In Design Edit task, where models must modify existing code according to user instructions, difficulty correlates with the complexity and scope of the modification instructions. We adopt the annotator-provided difficulty labels from Section IV-C as our ground truth difficulty classification.
@@ VI-C RQ3: How does varying task difficulty impact MLLM performance?
Table V shows the CLIP score for Design Generation task and MLLM score for Design Edit and Design Repair task under different difficulty levels.
Table V reveals distinct difficulty-related performance patterns across the three design tasks. In Design Generation, where difficulty stems from image complexity and size, top models show moderate degradation from  0.79-0.83 (Easy) to  0.69-0.73 (Hard). Design Edit tasks, with difficulty determined by the number of required operations, exhibit more pronounced drops, particularly for complex multi-operation scenarios where top models decline from 8.64-9.05 to 7.51-8.19. Design Repair shows the most severe degradation, where UI issue severity and required code modifications cause performance to plummet from 6.93-7.21 (Easy) to 4.19-6.93 (Hard), with smaller models experiencing catastrophic failures. These patterns indicate that visual complexity moderately affects generation, operational complexity significantly impacts editing, while code-level debugging presents the steepest performance barriers for current MLLMs.
@@
VI-D RQ4: To what extent do different input contexts affect MLLMs’ performance?
For the Design Edit and Design Repair tasks, we explore the impact of input images and codes on the results. Table VI shows the MLLM score on in Design Edit and Design Repair tasks under different context inputs. The results reveal distinct patterns regarding the utility of visual versus code information across different models and tasks.
@@ VI-D RQ4: To what extent do different input contexts affect MLLMs’ performance?
TABLE VI: Performance under different context inputs. Both denotes combining image and code.
@@
Figure 7: Compilation error distribution.
VI-E RQ5: What are the limitations of MLLMs when developing framework-based webpages?
VI-E1 Compilation Errors
In React development, the top three errors are “Unexpected Token”, “Expression Expected”, and “Use Client Missing”, with Qwen-7B producing the highest number of compilation errors (95 total) while GPT-4o demonstrates superior performance with only 3 total errors.
For Vue development, “Missing End Tag”, “Unexpected EOF”, and “Attribute Error” are dominated errors. Qwen-7B shows the most compilation issues (102), while Claude-3.7 achieves near-perfect compilation success with just 1 error.
Angular development presents a different error profile, with “Incomplete Block”, “Component Import Error”, and “Component Export Error” as the primary issues. Qwen-7B continues to struggle with 73 total errors, while Pixtral-124B shows improved performance compared to other frameworks.
@@ VI-E1 Compilation Errors
To verify MLLM’s ability to solve compilation errors, we sample 30 webpages with diverse compilation errors and prompt MLLMs to fix them. The results are shown in the Table VIII. The overall average repair rate across all models and frameworks is 0.53, indicating that MLLMs still face challenges in fixing front-end errors.
TABLE VII: Compile error fix ratio of MLLMs.
@@
TABLE VIII: Ratio of webpages using component-based design.
@@
VI-E2 Component-based Implementation Limitation
Component-based design is a method of breaking down user interfaces into reusable, self-contained parts called components. This approach improves efficiency and scalability in website development. Table VIII shows the proportion of webpages implemented by MLLMs using component-based design across three frameworks. The result reveals that MLLMs demonstrate remarkably low adoption rates of component-based implementation, with average 0.24%, 5% and 19% rate on React, Vue and Angular, respectively.

### v2

@@ 5.4. Metric
Visual Metrics. (1) CLIP (Radford et al., 2021) is applied to measure the semantic similarity between the generated and original webpages. (2) SSIM (Wang et al., 2004) (Structural Similarity Index Measure) score is applied to calculate the structure similarity. It evaluates the layout and compositional accuracy, emphasizing the spatial arrangement and structural similarities between the generated and original webpages.
@@
Code Metrics. (1) Compilation Success Rate (CSR). This metric represents the percentage of generated code that compiles successfully without errors. Assume that the total number of samples is N and the number of samples compiled successfully is S, then $CSR=\frac{S}{N}$. (2) Code Modification Location Similarity (CMLS). We employ the Jaccard similarity (Thada and Jaglan, 2013) to quantify the precision of code modifications on design edit and design repair tasks by comparing the sets of code modification operations between the ground truth and generated code. We first parse the code into an Abstract Syntax Tree (AST), and then identify the operation (i.e., add, delete and modify) by the AST node comparison. Let $A$ represent the set of code modification operations in the ground truth code and $B$ represent the set of code modification operations in the generated code. The CMLS is formally defined as: $CMLS(A,B)=\frac{|A\cap B|}{|A\cup B|}$. (3) Code Modification Content Similarity (CMCS). While CMLS focuses on modification locations, CMCS evaluates the semantic correctness of modified code content. Given modification operation sets $A$ and $B$, we match operations based on AST node correspondence and denote the matched pairs as $M\subseteq A\times B$. For each $(a,b)\in M$, CodeBLEU (ren2020codebleu) is used to compute content similarity, while unmatched operations receive a score of zero. The CMCS is defined as: $CMCS(A,B)=\frac{1}{|A\cup B|}\sum_{(a,b)\in M}\text{CodeBLEU}(a,b)$.
@@
MLLM-as-Judge Metrics. MLLMs have shown great performance in assisting judges across diverse modalities (Chen et al., 2024; Wang et al., 2025). Therefore, we prompt GPT-4o (OpenAI, 2024) to determine whether the model meets the user’s requirements on the design edit task and resolve the design issues on the design repair task, and output an MLLM score between 0 and 10 with detailed explanations (0-3 denotes the poor edit/repair, 4-6 denotes partial edit/repair, 7-8 denotes Good edit/repair and 9-10 denotes excellent edit/repair). The detailed judge prompts can be found in our repo [^3] https://github.com/WebPAI/DesignBench/blob/main/code/evaluator/metric.py. We employ the same five PhD students to conduct human validation studies to verify the reliability and consistency of our MLLM-as-Judge approach. Final validation are determined by majority voting. The Kappa score for inter-annotator agreement is 0.8648 and 0.8428 on design edit and repair task respectively, indicating ”almost perfect” consensus. For design edit task, we sample 359 samples and validate this MLLM score through human evaluation, which can achieve an average accuracy of 95.54%. For design repair task, we sample 111 samples and validate this MLLM score through human evaluation, which can achieve an average accuracy of 91.89%.
Table 4. Studied Multimodal Large Language Models.
@@
Pixtral (Agrawal et al., 2024) Pixtral-12B-2409 Pixtral-12B 12B
Pixtral (Agrawal et al., 2024) Pixtral-large-latest Pixtral-124B 124B
Qwen (Bai et al., 2025) Qwen2.5-VL-7B-Instruct Qwen-7B 7B
Qwen (Bai et al., 2025) Qwen2.5-VL-72B-Instruct Qwen-72B 72B
LLama (AI, 2025) Llama-3.2-11B-Vision-Instruct Llama-11B 11B
LLama (AI, 2025) Llama-3.2-90B-Vision-Instruct LLama-90B 90B
Gemini (Google, 2024) Gemini-2.0-Flash Gemini-2.0 -
GPT (OpenAI, 2024) GPT-4o-2024-11-20 GPT-4o -
Claude (Anthropic, 2024) Claude-3-7-sonnet-20250219 Claude-3.7 -
5.5. Research Questions
@@
  • RQ3: (Difficulty Influence) How does varying task difficulty impact MLLM performance?
  • RQ4: (Context Influence) To what extent do different input contexts affect MLLMs’ performance?
@@
  • RQ6: (Failure analysis) What mistakes do MLLMs make on the three tasks?
6. Experiment Results
Table 5. The model performance on DesignBench under different tasks and different frameworks. Bold numbers on a dark red background indicate the maximum values, and an underline with a light red background denotes the second-best value.
@@ 6. Experiment Results
SSIM React 0.6912 0.6655 0.6577 0.6463 0.6255 0.6478 0.5576 0.6774 0.0832
Vue 0.7155 0.6809 0.6136 0.4228 0.2727 0.6339 0.6432 0.5681 0.0411
Angular 0.4974 0.5196 0.5251 0.4883 0.4930 0.5138 0.3996 0.4321 0.0814
Vanilla 0.7004 0.6788 0.6820 0.6346 0.6687 0.6471 0.6229 0.6673 0.6476
CSR React 0.9541 0.9725 0.9083 0.9450 0.8991 0.9725 0.8532 0.9541 0.1284
@@
CSR React 1.0000 0.9815 1.0000 0.9167 0.7963 0.9907 1.0000 0.9907 0.4815
@@
CMLS React 0.5555 0.6159 0.5325 0.3836 0.3272 0.5306 0.5209 0.5930 0.1890
Vue 0.5961 0.5090 0.5414 0.1924 0.1443 0.5290 0.3990 0.4720 0.1705
Angular 0.7696 0.7267 0.6838 0.5057 0.4041 0.5176 0.5933 0.7016 0.3453
Vanilla 0.5862 0.5727 0.5645 0.3865 0.3113 0.5644 0.4764 0.5782 0.3125
CMCS React 0.4551 0.5122 0.4058 0.3022 0.2425 0.4149 0.4189 0.4872 0.1319
Vue 0.4807 4232 0.4292 0.1521 0.1016 0.4168 0.3199 0.3837 0.1139
Angular 0.7080 0.6442 0.5945 0.4042 0.3203 0.4075 0.4790 0.6225 0.2240
Vanilla 0.5357 0.5335 0.5177 0.3532 0.2757 0.5260 0.4311 0.5395 0.2543
@@
CSR React 1.0000 1.0000 1.0000 0.9286 0.9643 1.0000 0.9643 0.9286 0.2857
@@
CMLS React 0.6340 0.4833 0.3536 0.1389 0.0666 0.4551 0.1959 0.3616 0.0784
Vue 0.5089 0.4865 0.3489 0.1533 0.1494 0.4422 0.3931 0.2451 0.0635
Angular 0.6759 0.5566 0.4785 0.3876 0.2729 0.5086 0.4995 0.6414 0.0898
Vanilla 0.6088 0.6659 0.4867 0.4237 0.2759 0.5563 0.5447 0.4990 0.1503
CMCS React 0.5420 0.3872 0.2548 0.0881 0.0365 0.3411 0.1331 0.2567 0.0563
Vue 0.3924 0.3748 0.2880 0.1024 0.0912 0.3377 0.3244 0.1800 0.0349
Angular 0.5706 0.4527 0.3617 0.2711 0.1463 0.3894 0.3606 0.5773 0.0368
Vanilla 0.5859 0.6402 0.4578 0.3780 0.2158 0.5404 0.5279 0.4780 0.1172
6.1. RQ1: How do MLLMs perform across distinct front-end tasks?
Table 5 presents the performance of nine MLLMs on three tasks in three front-end frameworks React, Vue, and Angular. Vanilla denotes the webpage developed by vanilla HTML/CSS.
@@ 6.1. RQ1: How do MLLMs perform across distinct front-end tasks?
In Design Generation tasks, MLLMs face two primary bottlenecks: compilation errors and visual inaccuracies. The compilation rates reveal significant framework-dependent challenges, with Angular showing the lowest success rates (0.6867-0.7590 for top models) compared to React and Vue (>>0.83 for top models). Additionally, the moderate CLIP scores (0.5963-0.6324 even for best performers) indicate substantial room for improvement in generating visually accurate webpage layouts. The performance gap between vanilla HTML (highest CLIP scores) and framework-based implementations further suggests that the syntactic complexity in modern frameworks exacerbates challenges in both compilation and visual rendering.
In the Design Edit and Design Repair tasks, the primary bottleneck for MLLMs lies in accurately localizing and modifying the code segments that require changes. This limitation is reflected by the substantially lower CMLS and CMCS scores compared to compilation success rates across all models. In the Design Edit task, even top-performing models such as Claude-3.7 achieve CMLS scores of only 0.5555–0.7696 and CMCS scores of 0.4551–0.7080. A similar trend is observed in the Design Repair task, where Claude-3.7 attains CMLS values ranging from 0.5089 to 0.6759 and CMCS values from 0.3924 to 0.5859. These results indicate that, despite generating code that successfully compiles, MLLMs still struggle to precisely identify and semantically modify the target code regions.
Finding 3: MLLMs exhibit task-specific bottlenecks: Design Generation is challenged by compilation errors and visual inaccuracies, while Design Edit and Repair are mainly limited by deficiencies in code localization and modification.
6.2. RQ2: What is the comparative performance of MLLMs when applied to different development frameworks?
@@
Figure 6. Performance of webpages implemented by different frameworks. y-axis denotes the framework used to actually implement the webpage and x-axis represents the framework used by the model.
@@
6.3. RQ3: How does varying task difficulty impact MLLM performance?
Table 6. Performance under different difficulty levels.
Model Design Generation Design Edit Design Repair
Easy Medium Hard Easy Medium Hard Easy Medium Hard
Claude-3.7 0.86 0.77 0.51 8.64 8.32 7.61 7.21 6.81 6.93
GPT-4o 0.84 0.74 0.43 8.73 7.95 8.19 6.93 6.46 5.53
Gemini-2.0 0.84 0.74 0.743 9.05 7.84 7.51 7.07 6.21 4.19
Llama-90B 0.69 0.65 0.47 6.54 6.83 5.68 5.35 4.90 4.39
Llama-11B 0.68 0.62 0.46 5.88 5.05 5.24 3.91 4.07 2.41
Pixtral-124B 0.78 0.73 0.53 8.69 7.91 6.98 7.30 6.56 5.66
Pixtral-12B 0.67 0.68 0.50 8.26 7.63 6.71 6.76 5.84 5.30
Qwen-72B 0.83 0.71 0.44 7.78 6.81 6.72 7.12 6.06 5.57
Qwen-7B 0.56 0.53 0.32 3.46 2.83 2.71 1.09 1.63 1.05
We categorize samples into different difficulty levels to systematically evaluate MLLMs’ performance across varying complexity scenarios.
For Design Generation task, which requires generating code from scratch, we calculate the complexity based on the image size $I$, UI elements number $U$, color variety $C$ (i.e., the number of colors in the UI) and layout complexity $L$ (i.e., number of layout methods + number of nesting levels + number of CSS layout attributes). We then normalize these scores to between 0 and 100. The complexity score: $S=0.25\times I+0.25\times U+0.25\times C+0.25\times L$. Scores greater than 80 are complex, less than 30 are simple, and between 30 and 80 are medium.
In Design Edit task, where models must modify existing code according to user instructions, difficulty correlates with the complexity and scope of the modification instructions. We adopt the annotator-provided difficulty labels from Section 4.3 as our ground truth difficulty classification.
@@ 6.3. RQ3: How does varying task difficulty impact MLLM performance?
Table 6 shows the CLIP score for Design Generation task and MLLM score for Design Edit and Design Repair task under different difficulty levels.
Table 6 reveals distinct difficulty-related performance patterns across the three design tasks. In Design Generation, where difficulty stems from image complexity and size, top models show moderate degradation from  0.79-0.83 (Easy) to  0.69-0.73 (Hard). Design Edit tasks, with difficulty determined by the number of required operations, exhibit more pronounced drops, particularly for complex multi-operation scenarios where top models decline from 8.64-9.05 to 7.51-8.19. Design Repair shows the most severe degradation, where UI issue severity and required code modifications cause performance to plummet from 6.93-7.21 (Easy) to 4.19-6.93 (Hard), with smaller models experiencing catastrophic failures. These patterns indicate that visual complexity moderately affects generation, operational complexity significantly impacts editing, while code-level debugging presents the steepest performance barriers for current MLLMs.
@@
6.4. RQ4: To what extent do different input contexts affect MLLMs’ performance?
For the Design Edit and Design Repair tasks, we explore the impact of input images and codes on the results. Table 7 shows the MLLM score on in Design Edit and Design Repair tasks under different context inputs. The results reveal distinct patterns regarding the utility of visual versus code information across different models and tasks.
@@ 6.4. RQ4: To what extent do different input contexts affect MLLMs’ performance?
Table 7. Performance under different context inputs. Both denote combining image and code.
@@
6.5. RQ5: The limitations of MLLMs when developing framework-based webpages?
6.5.1. Compilation Errors
In React development, the top three errors are “Unexpected Token”, “Expression Expected”, and “Use Client Missing”, with Qwen-7B producing the highest number of compilation errors (95 total) while GPT-4o demonstrates superior performance with only 3 total errors. For Vue development, “Missing End Tag”, “Unexpected EOF”, and “Attribute Error” are dominated errors. Qwen-7B shows the most compilation issues (102), while Claude-3.7 achieves near-perfect compilation success with just 1 error. Angular development presents a different error profile, with “Incomplete Block”, “Component Import Error”, and “Component Export Error” as the primary issues. Qwen-7B continues to struggle with 73 total errors, while Pixtral-124B shows the best performance.
@@ 6.5.1. Compilation Errors
Figure 7. Compilation error distribution.
@@
To verify MLLM’s ability to solve compilation errors, we sample 30 webpages with diverse compilation errors and prompt MLLMs to fix them. The results are shown in the Table 9. The overall average repair rate across all models and frameworks is 0.53, indicating that MLLMs still face challenges in fixing front-end errors.
Table 8. Compile error fix ratio.
@@
Table 9. Ratio of using component-based design.
@@
6.5.2. Component-based Implementation Limitation
Component-based design is a method of breaking down user interfaces into reusable, self-contained parts called components. This approach improves efficiency and scalability in website development. Table 9 shows the proportion of webpages implemented by MLLMs using component-based design across three frameworks. The result reveals that MLLMs demonstrate remarkably low adoption rates of component-based implementation, with average 0.24%, 5% and 19% rate on React, Vue and Angular, respectively.